OpenStudy (anonymous):

wells fargo bank loaned 30,000, some at an annual rate of 7% and some at an annual rate of 3%, if the income of these loans was 1,380 how much was loaned at 7% and how much was loaned at 3%?

OpenStudy (gw2011):

You need to set up two equations: Let x=loaned at 7% Let y=loaned at 3% x+y=30,000 0.07x+0.03y=1380 Let x=30,000-y and substitute: 0.07(30,000-y)+0.03y=1380 2100-0.07y+0.03y=1380 2100-0.04y=1380 2100-1380=0.04y rearranging this you get: 0.04y=2100-1380 0.04y=720 y=18,000 substituting this into one of the equations: x+y=30,000 x+18,000=30,000 x=30,000-18,000 x=12,000 Therefore, 12,000 was loaned at 7% and 18,000 was loaned at3%